Using the appropriate interest table, answer the following questions. (Each case is independent of the others.) Your answer iplease provide me explaination .
Answer #1

solution 1:

Future value = $4,000 * cumulative FV factor at 8% for 20 periods of annuity due

= $4,000 * 49.42292 = $197,692

Solution 2:

Present value = $2,500 * Cumulative PV factor at 5% for 30 periods of annuity due

= $2,500 * 16.14107 = $40,353

Solution 3:

Future value = $2,000 * cumulative FV factor at 10% for 15 periods of annuity due

= $2,000 * 34.94973 = $69,899

Solution 4:

Present value = $1,000 * Cumulative PV factor at 9% for 6 periods of annuity due

= $1,000 * 4.88965 = $4,890